A population consists of the numbers 1, 2, 3, and 4 with the sample size of 2 (without replacement). Which mean that has a probability of 2/6 or 1/3 or 0.33?




1
Expert's answer
2022-04-18T16:30:41-0400

The number of possible samples which can be selected without replacement is

"\begin{pmatrix}
 N \\
 n
\end{pmatrix}=\cfrac{N! } {n! \cdot(N-n)! }=\cfrac{4! } {2! \cdot2! }=6."



"\textnormal{Sample\qquad\qquad Mean}\\ 
\ \\
1,2\qquad \cfrac{1+2}{2}=1.5\\
1,3\qquad \cfrac{1+3}{2}=2\\
1,4\qquad \cfrac{1+4}{2}=2.5\\
2,3\qquad \cfrac{2+3}{2.5}=2.5\\
2,4\qquad \cfrac{2+4}{2}=3\\
3,4\qquad \cfrac{3+4}{2}=3.5."


We see that the mean = 2.5 met twice, its probability is "\cfrac{2}{6}=\cfrac{1}{3}."

All the other means met once, their probabilities are "\cfrac{1}{6}."